# A High-Performance Evolutionary Multiobjective Community Detection Algorithm

### v1

@@ 5 Conclusion and Future Work
We introduced HP-MOCD, a high-performance evolutionary multi-objective algorithm for large-scale network community detection. Built on the NSGA-II framework, HP-MOCD incorporates topology-aware crossover and mutation operators, coupled with a parallelized, efficient implementation, ensuring scalability and practical applicability. Unlike traditional single-objective heuristics and earlier MOEA-based approaches, HP-MOCD provides a principled mechanism to capture diverse structural trade-offs via Pareto-optimal partitions. The algorithm consistently outperformed other evolutionary methods in runtime and scalability, maintaining competitive or superior accuracy, even in complex or noisy networks.
These results significantly advance the state-of-the-art in multi-objective community detection. While most existing methods are limited by high computational costs or inflexibility, HP-MOCD demonstrates multi-objective expressiveness combined with high-performance execution. Furthermore, its ability to yield diverse, high-quality partitions makes the algorithm especially suitable for exploratory network analysis and downstream tasks benefiting from structural alternatives.
Future work can extend the algorithm to support overlapping communities and more general network types, such as weighted, directed, or temporal graphs. Another promising path involves improving robustness under high mixing regimes ($\mu\to 1$), potentially through hybrid strategies or advanced selection mechanisms. Finally, we intend to explore integration with NSGA-III [33], and other recent MOEA frameworks to further improve performance in scenarios involving more than two objectives, such as modularity, conductance, or community balance.
@@ Appendix A Space Complexity Analysis
Let $|V|$ represent the number of nodes, $|E|$ the number of edges, $N_{p}$ the population size, and $d_{\max}=\max_{v\in V}d(v)$ the maximum node degree. The graph $G$ itself is stored once, requiring $O\bigl{(}|V|+|E|\bigr{)}$ space. Each individual in the population is a labeling of all $|V|$ nodes, thus one individual occupies $O\bigl{(}|V|\bigr{)}$ space. At any generation, HP-MOCD holds both the parent population $P_{t}$ and the offspring population $Q_{t}$. This amounts to $2N_{p}$ individuals, consuming
@@ Appendix B Role of Multi-Threaded Architecture in Algorithm Performance
To quantify the impact of multi-threading, we benchmarked the algorithm on a high-performance system featuring an AMD Ryzen Threadripper 3960X processor (24 physical cores, 48 logical threads, @ 3.70GHz) and 128GB DDR4 RAM. The number of execution threads was systematically varied from 1 to 48. These tests utilized the same synthetic graph datasets with $n\in[10,000,30,000,50,000]$ nodes as detailed in our scalability analysis (Section 4.1.1). Each configuration was executed five times, a sufficient number given the consistently low standard deviation observed in execution times.
@@ Appendix C Selection of Genetic Algorithm Parameters
Aggregate Quality Metric
@@
which balances the two scores and penalizes configurations more strongly when either AMI or NMI is low. While AMI and NMI are widely used, the methodology remains compatible with other external or internal evaluation metrics, depending on the application context.
(a) H versus E
(b) Execution Time versus E
Figure 8: Harmonic mean of AMI and NMI versus GA effort $E=P\times G$, and execution time versus GA effort $E=P\times G$, for different graph sizes.
GA Effort Metric
@@
Figure 8(a) and Figure 8(b) display, respectively, the harmonic-mean quality and runtime as functions of $E$, with separate curves for each $n$.

### v2

@@ 6 Conclusion and Future Work
We introduced HP-MOCD, a high-performance evolutionary multi-objective algorithm for large-scale network community detection. Built upon the NSGA-II framework, HP-MOCD incorporates topology-aware crossover and mutation operators, coupled with a parallelized and efficient implementation that ensures practical scalability. In contrast to traditional single-objective heuristics and earlier MOEA-based methods, HP-MOCD produces a Pareto front of diverse, high-quality partitions, enabling a principled exploration of trade-offs between structural cohesion and ground-truth alignment.
Extensive experiments on synthetic and real-world datasets show that HP-MOCD consistently outperforms prior evolutionary methods in both runtime and robustness, while delivering competitive accuracy, particularly in scenarios involving noisy, ambiguous, or large-scale networks. Even in cases where single-objective methods such as Louvain or Leiden yield higher scores in individual metrics, HP-MOCD maintains a more balanced performance profile across AMI, NMI, modularity, and F1-score, highlighting the benefits of its multi-objective formulation.
Future work can extend the algorithm to support overlapping communities and more general network types, such as weighted, directed, or temporal graphs. Another promising path involves improving robustness under high mixing regimes ($\mu\to 1$), potentially through hybrid strategies or advanced selection mechanisms. An additional direction includes investigating the resilience of communities under node or edge failures, as explored in prior work on failure-resilient communities [48, 49]. We also envision extending our experimental evaluation by incorporating additional datasets and more baseline algorithms, such as those proposed by [Ghoshal:2019, Ghoshal:2021], to more thoroughly assess the strengths and limitations of HP-MOCD in diverse scenarios. Furthermore, accurate and scalable community detection can serve as a foundation for downstream applications such as diffusion source inference and diffusion containment in complex networks, which increasingly rely on community structure to enhance performance [50, 51, 52]. Lastly, we plan to explore integration with NSGA-III [53] and other recent MOEA frameworks to support higher-dimensional objectives, such as conductance and community balance, thereby expanding the scope of multi-objective community detection.
@@ Appendix A Space Complexity Analysis
[…]
@@ Appendix B Role of Multi-Threaded Architecture in Algorithm Performance
To quantify the impact of multi-threading, we benchmarked the algorithm on a high-performance system featuring an AMD Ryzen Threadripper 3960X processor (24 physical cores, 48 logical threads, @ 3.70GHz) and 128GB DDR4 RAM. The number of execution threads was systematically varied from 1 to 48. These tests utilized the same synthetic graph datasets with $n\in[10,000,30,000,50,000]$ nodes as detailed in our scalability analysis (Appendix C). Each configuration was executed five times, a sufficient number given the consistently low standard deviation observed in execution times.
@@ Appendix C Selection of Genetic Algorithm Parameters
C.1 Aggregate Quality Metric
@@
which balances the two scores and more strongly penalizes configurations where either AMI or NMI is low. While AMI and NMI are widely used, the methodology remains compatible with other external or internal evaluation metrics, depending on application context.
C.2 GA Effort Metric
@@
Figure 9(a) and Figure 9(b) display, respectively, the harmonic mean quality and runtime as functions of $E$, with separate curves for each $n$.
(a) H versus E
(b) Execution Time versus E
Figure 9: Harmonic mean of AMI and NMI versus GA effort $E$, and execution time versus GA effort $E$, for different graph sizes.
@@ C.3 Chosen Configuration
Appendix D Crossover and Mutation Example
Figure 10 provides a visual example of the crossover and mutation operators described in Section 4.3. These examples help illustrate how the genetic variation mechanisms alter individuals during the evolutionary process.
1234567812463578(a) Parent 1 (selected)1234567812354678(b) Parent 2 (selected)1234567812384567(c) Parent 3 (selected)12345678123456784 (d) Crossover Offspring 12345678123456784 Neighbors of 4 are {1,2,3,5}. Communities are {B,B,B,G}. Most frequent is Blue. (e) Mutated Offspring
Figure 10: An illustrative example of the Crossover and Mutation operators. (a), (b) and (c) show three parent solutions selected. (d) The offspring is created by a majority vote for each node’s community assignment, as determined by Equation 10. The result is a novel partition, different from all parents. (e) The mutation operator inspects the neighborhood of node 4, finds that blue is the most frequent community, and flips its color to perform a final correction, recovering the optimal partition.